OpenStudy (anonymous):

(Watch nobody come to answer, even when I mention people I fanned for helping me)

OpenStudy (anonymous):

What is the solution to 3x^2 + 4x + 8 = 0? A) x = -4 ± 4√5 B) x = 2 ± 2√5 / 3 C) x = -4 ± 4√5 / 6 D) x = -2 ± 2i√5 / 3

Arcadiouse (arcadiouse):

Sure. I will help.

Arcadiouse (arcadiouse):

Oh sorry, this is out of my range.

OpenStudy (anonymous):

@itrout can you get lost ???

OpenStudy (anonymous):

I only need 1 person to help

OpenStudy (littlet):

@ganeshie8

Arcadiouse (arcadiouse):

@Qwertty123

OpenStudy (anonymous):

And I don't appreciate your little comment earlier. Why did you have to pull sexism into the argument?

OpenStudy (anonymous):

Oh you know @Qwertty123 ?

Arcadiouse (arcadiouse):

Yas

OpenStudy (anonymous):

I don't tag mod or ambassadors btw

OpenStudy (anonymous):

Qwertty is my bff

Arcadiouse (arcadiouse):

Oh your telling me the rules now.

OpenStudy (anonymous):

HEY QWERTTS WE TALKIN BOUT YU

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

i am new here and cute who wants to help<3

OpenStudy (qwertty123):

Hey! :D Sorry I only got notifications now..

OpenStudy (anonymous):

GTFO CHECK

OpenStudy (qwertty123):

:o @checkm8 I know you! Why are you on a alt?

OpenStudy (anonymous):

HAHAHAHAH

OpenStudy (anonymous):

?

OpenStudy (anonymous):

CAN I GET SOME MATH HELP LIKE GEEZ 9TH GRADE IS TOUGH

OpenStudy (qwertty123):

Okay.. so do you know how to solve 3x^2 + 4x + 8 = 0?

OpenStudy (qwertty123):

or are you clueless..

Arcadiouse (arcadiouse):

Everybody is stalking this lmao

OpenStudy (anonymous):

tell us \(\Huge YOUR\) attempt

OpenStudy (qwertty123):

Brownie.. can you please tell me what you have so far?

OpenStudy (qwertty123):

Step 1 : Equation at the end of step 1 : (3x2 + 4x) - 8 = 0 Trying to factor by splitting the middle term 2.1 Factoring 3x2+4x-8 The first term is, 3x2 its coefficient is 3 . The middle term is, +4x its coefficient is 4 . The last term, "the constant", is -8 Step-1 : Multiply the coefficient of the first term by the constant 3 • -8 = ?

MsBrains (ms-brains):

Great job helping this user! @Qwertty123

OpenStudy (qwertty123):

Thanks so much! :D I am just wait for this person to help me out...

OpenStudy (cheesecakekitten):

i THINK she got kicked lol

OpenStudy (qwertty123):

Back! but my connection isn't great... I might leave again.. anyways what is.. 3 • -8 = ?

OpenStudy (cheesecakekitten):

@BrownieCat r u dere?????

OpenStudy (cheesecakekitten):

sis ;-;

OpenStudy (fortytherapper):

"(Watch nobody come to answer, even when I mention people I fanned for helping me)" Form of reverse physiology?

OpenStudy (cheesecakekitten):

shes gone qaq

OpenStudy (qwertty123):

o.O What does that have to do with this math.. XD

OpenStudy (cheesecakekitten):

qwertt bertt lol

OpenStudy (cheesecakekitten):

qwatt batt

OpenStudy (cheesecakekitten):

qwitt bitt

OpenStudy (cheesecakekitten):

qwutt butt

OpenStudy (cheesecakekitten):

lol

OpenStudy (qwertty123):

o.O Excuse me!

alones (alones):

@CheesecakeKitten I though you were very nice young lady ;3 turns out ...

OpenStudy (cheesecakekitten):

QWYTT BYTT

OpenStudy (qwertty123):

@BrownieCat can you please help me?

OpenStudy (cheesecakekitten):

i THINK SHE WAS BANNED

OpenStudy (cheesecakekitten):

Wut

alones (alones):

Kityten please don't spam;~;

OpenStudy (qwertty123):

Would you stop spamming..

OpenStudy (cheesecakekitten):

srry

OpenStudy (cheesecakekitten):

What is the solution to 3x^2 + 4x + 8 = 0? A) x = -4 ± 4√5 B) x = 2 ± 2√5 / 3 C) x = -4 ± 4√5 / 6 D) x = -2 ± 2i√5 / 3

OpenStudy (qwertty123):

I am trying to help brownie...

alones (alones):

BROWNIE MAKE A NEW POST I'LL TYR TO HELP TOO ;)